OpenStudy (anonymous):

how do you graph (y-2)^2=-2(x+2)

OpenStudy (anonymous):

Parabola

OpenStudy (anonymous):

What is the foci of this equation? And shouldn't it be opening to the left?

OpenStudy (anonymous):

@Directrix

sam (.sam.):

generally for any given equation like (y + h)^2 = (x + k). Vertex = (-h,-k) or you could use complete the square for y \[(y-2)^2=-2 (x+2)\] \[-2 (x+2)=(y-2)^2\] \[x=-\frac{1}{2} (y-2)^2-2\] Vertex (-2,2) for focus, the coefficient of the unsquared part is -2, 4p=-2, p=-1/2 focus=-1/2 + (-2)=-5/2 Focus (-5/2,2)

OpenStudy (anonymous):

Do you mind if I ask another one on here? @.Sam.

sam (.sam.):

yes?

OpenStudy (anonymous):

y^2+6y+8x-7=0 How do you graph this parabola?

sam (.sam.):

arrange it like this, \[x=-\frac{y^2}{8}-\frac{3 y}{4}+\frac{7}{8}\] complete the square \[x= -\frac{1}{8} (y+3)^2+2\] can you find the vertex and foci from here?

OpenStudy (anonymous):

Well I completed the square instead of doing it that way. But how do you graph it?

sam (.sam.):

you got the vertex (2,-3), and foci, (0,-3)